LunaBechtold:

Simplify 5 to the power of 2x5 to the power of 4x5-4x5-2.

P3achPounder:

@hyper

P3achPounder:

#ap

P3achPounder:

@astroboy

Syrnn:

real

jhonyy9:

\[5^{2} × 5^{4} × 5 - 4 × 5 -2 = ?\] do you think it hence ?

Syrnn:

@jhonyy9 wrote:
\[5^{2} × 5^{4} × 5 - 4 × 5 -2 = ?\] do you think it hence ?
thatd be 25 x 625 x 5 = 78,125 and then 78,125 - 4 = 78,121 x 5 = 390,605 - 2 = 390,603 😜

Syrnn:

@syrnn wrote:
@jhonyy9 wrote:
\[5^{2} × 5^{4} × 5 - 4 × 5 -2 = ?\] do you think it hence ?
thatd be 25 x 625 x 5 = 78,125 and then 78,125 - 4 = 78,121 x 5 = 390,605 - 2 = 390,603 😜
no no no i got it wrong oh boy its 15603 maybe idk man i used a calculator to look like a big boy

Spectrum:

Use P.E.M.D.A.S. Easy peasy.